OpenStudy (anonymous):

Amistre64 can you help me?

OpenStudy (anonymous):

\[\int\limits_{}^{}15/2(t+1)^{1/2}\]

OpenStudy (anonymous):

we can do this

OpenStudy (anonymous):

\[\int\frac{15}{2\sqrt{t+1}}dt\]?

OpenStudy (anonymous):

can someone help me?

OpenStudy (anonymous):

is that the question? first off the constants come right out front of the integral sign

OpenStudy (anonymous):

yes that is the question and thank you

OpenStudy (anonymous):

\[\frac{15}{2}\int \frac{1}{(t+1)^{\frac{1}{2}}}\]

OpenStudy (anonymous):

now a very simple u - substitution, not even really necessary. put \[u=t+1\] and \[du=dt\] so you get \[\frac{15}{2}\int u^{-\frac{1}{2}}du\]

OpenStudy (anonymous):

now use power rule backwards to get the anti - derivative

OpenStudy (anonymous):

\[\int u^{-\frac{1}{2}}=2u^{\frac{1}{2}}\]

OpenStudy (anonymous):

hope that is clear. just add one to the exponent and then divide by that exponent.

OpenStudy (anonymous):

yesthat is clear

OpenStudy (anonymous):

now replace u by t +1 and remember there is a fraction out front to get \[\frac{15}{2}\times 2 (t+1)^\frac{1}{2}=15\sqrt{t+1}\]

OpenStudy (anonymous):

and i guess you need that stupid +C at the end

OpenStudy (anonymous):

easy check is to take the derivative and see that you have it correct.

OpenStudy (anonymous):

yes you need that stupid C

OpenStudy (anonymous):

ok hope all steps are clear. little practice with ones like this you do the "u substitution" in your head since in this case du = dx

OpenStudy (anonymous):

btw ictrees sent me to you. she is incognito